1+2*(3-4)^2 =
-5*(-2+4)^2(2^3/4) =

thnaks in advance :)

Answer by rfer(16322)

1+2*(3-4)^2=
1+2*1=3
------------
-5*(-2+4)^2(2^3/4)=
-5*(4)(1.68)=-33.6